OpenStudy (anonymous):

Let f(x)=-2x-7 and g(x)=-4x+6 find (g (open o) f) (-5)

OpenStudy (anonymous):

this time first we do f, then we do g: \[f(-5)=-2\times (-5)-7=10-7=3\]

OpenStudy (anonymous):

\[g(3)=-4\times 3 +6=-12+6=-6\]